Solve the following system of linear equations by the elementary row operations: 5

𝑝 + 2π‘ž βˆ’ 2π‘Ÿ βˆ’ 𝑑 = 0

2𝑝 + 5π‘ž βˆ’ 3π‘Ÿ βˆ’ 𝑑 = 1

3𝑝 + 8π‘ž βˆ’ 4π‘Ÿ βˆ’ 𝑑 = 2

𝑝 + 5π‘ž + π‘Ÿ + 2𝑑 = 3


1
Expert's answer
2021-03-09T00:20:54-0500

Solve the following system of linear equations by the elementary row operations:

5𝑝 + 2π‘ž βˆ’ 2π‘Ÿ βˆ’ 𝑑 = 0

2𝑝 + 5π‘ž βˆ’ 3π‘Ÿ βˆ’ 𝑑 = 1

3𝑝 + 8π‘ž βˆ’ 4π‘Ÿ βˆ’ 𝑑 = 2

𝑝 + 5π‘ž + π‘Ÿ + 2𝑑 = 3

Solution:

"𝑝 + 5𝑞 + 𝑟 + 2𝑡 = 3\\
5𝑝 + 2𝑞 − 2𝑟 − 𝑡 = 0\\
2𝑝 + 5𝑞 − 3𝑟 − 𝑡 = 1\\
3𝑝 + 8𝑞 − 4𝑟 − 𝑡 = 2"


Multiply 1-st equation by -5 and add it to the second equation

Multiply 1-st equation by -2 and add it to the third equation

Multiply 1-st equation by -3 and add it to the fourth equation


"𝑝 + 5𝑞 + 𝑟 + 2𝑡 = 3\\
0 -23𝑞 − 7𝑟 − 11𝑡 = -15\\
0 - 5𝑞 − 5𝑟 − 5𝑡 = -5\\
0 -7𝑞 − 7𝑟 −7 𝑡 = -7"


Multiply 2-nd equation by 23 and add it to the third equation


"𝑝 + 5𝑞 + 𝑟 + 2𝑡 = 3\\
0 +𝑞 +𝑟 +𝑡 = 1\\
0 -23𝑞 − 7𝑟 − 11𝑡 = -15\\
0 +0 +0 +0 = 0"


"𝑝 + 5𝑞 + 𝑟 + 2𝑡 = 3\\
0 +𝑞 +𝑟 +𝑡 = 1\\
0 +0 +16𝑟 +12𝑡 =8\\
0 +0 +0 +0 = 0"


"𝑝 + 5𝑞 + 𝑟 + 2𝑡 = 3\\
0 +𝑞 +𝑟 +𝑡 = 1\\
0 +0 +𝑟 +0.75𝑡 =0.5\\
0 +0 +0 +0 = 0"


Multiply third equation by -1 and add it to the first equation

Multiply third equation by -1 and add it to the second equation


"𝑝 + 5𝑞 + 0 + 1.25𝑡 = 2.5\\
0 +𝑞 +0 +0.25𝑡 = 0.5\\
0 +0 +𝑟 +0.75𝑡 =0.5\\
0 +0 +0 +0 = 0"


Multiply second equation by -5 and add it to the first equation


"𝑝 +0 + 0 + 0 = 0\\
0 +𝑞 +0 +0.25𝑡 = 0.5\\
0 +0 +𝑟 +0.75𝑡 =0.5\\
0 +0 +0 +0 = 0"


Answer:

"𝑝= 0\\
𝑞= 0.5-0.25𝑡\\
𝑟=0.5-0.75𝑡"